OpenStudy (anonymous):

Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

are the numbers after the letters supposed to be powers?

OpenStudy (anonymous):

I can't imagine it being anything else

OpenStudy (anonymous):

But do write your exponents with ^

jhonyy9 (jhonyy9):

so will be -9 -3*8 +9 +2*8 -9 = -24 +16 -9 = 16 -33 = -17 hope this will help you